OpenStudy (kittiwitti1):

#37 has already seen a few attempts to answer/help/assist but I’ve already got the answer. Please view the first two questions instead if you want to help. The questions are on work, energy and power. I apologize if I'm not able to answer right away because the site is down on my side - I can't connect to the question link in order to perform such an action. --- 48) Starting from rest, a 10.0 kg suitcase slides 3.00 m down a frictionless ramp inclined at 30.0 degrees from the floor. The suitcase then slides an additional 5.00 m along the floor before coming to a stop. Determine the following: A, the suitcase’s speed at the bottom of the ramp (I got that answer - it was 5.42 m/s) B, the coefficient of kinetic friction C, the change in mechanical energy due to friction 49) A light horizontal spring has a spring constant of 105 N/m. A 2.00 kg block is pressed against one end of the spring, compressing the spring 0.100 m. After the block is released, the block moves 0.250 m to the right before coming to rest. What is the coefficient of kinetic friction between the horizontal surface and the block? 37) A 215 g particle is released from rest at point A inside a smooth hemispherical bowl of radius 30.0 cm. Calculate the following: a, the gravitational potential energy at A relative to B b, the particle's kinetic energy at B c, the particle's speed at B d, the potential energy and kinetic energy at C

OpenStudy (kittiwitti1):

@AllTehMaffs

OpenStudy (anonymous):

hallo - it's all broke again!

OpenStudy (anonymous):

Picture?

OpenStudy (kittiwitti1):

actually i think i got it hold on

OpenStudy (anonymous):

very nice ^_^

OpenStudy (anonymous):

Maybe we do not need a picture, the potential energy differences will simply be m g times the height differences. Sliding (rather than rolling) means all the potential energy changes get transformed to kinetic energy changes, (1/2) m v^2. Total PE + KE will be constant, allowing calculation where we have both.

OpenStudy (kittiwitti1):

@douglaswinslowcooper I meant to type this earlier but OS wouldn't let me onto the question to do so... I've solved the equation.... @AllTehMaffs ON TO THE NEXT ONE!

OpenStudy (kittiwitti1):

@AllTehMaffs helpp D:

OpenStudy (anonymous):

Not sure which unanswered. You got velocity from loss of potential energy m g h = 0.5 m v^2 kinetic energy v=5.4 m/s Next note energy m g h is lost to friction work F d, where d = 5 m m g h = 147 J = 5 F, so F = 29.4 N mu = F/W, F is friction and W is weight mu = 29.4 N / 98 N = 0.3 coefficient of friction

OpenStudy (anonymous):

The change in mechanical energy was just m g h.

OpenStudy (anonymous):

Block and spring P.E = 0.5 k x^2 = 0.5 x 105 x .01 = 0.525 J or N-m Work = loss of that energy = force x distance 0.525 N-m= F x 0.25 m so F = .525/.250 = 2.1 N mu = F/weight = 2.1 / (2)(9.8) = 0.107.

OpenStudy (kittiwitti1):

@douglaswinslowcooper thanks for helping... but what question are you helping me with?

OpenStudy (anonymous):

@kittiwitti1 are these questions from MasteringPhysics?

OpenStudy (kittiwitti1):

@HomeschoolGrad idk?

OpenStudy (anonymous):

48) So yeah, you already have the velocity at the bottom of the ramp, so there are a few ways to do it. @douglaswinslowcooper 's way is super slick, and basically uses the answer to the last part of the question to find the force; since you know that the work is given by \[W = \Delta K\] where \[K_i = \frac{1}{2}mv^2 \rightarrow \text{where v is the v you've already found}\] \[K_f = 0 \rightarrow \text{since there is no velocity when the object is at rest!}\] and you can also note from conservation of energy from the top of the ramp to the bottom, \[K_i = U = mgh\] thus \[W = K_f - K_i = -mgh\] And you also know that \[W=Fd \cos \theta \rightarrow \text{d is distance over which the force acts} \\ \qquad \qquad \text{ and theta is the angle between the force and the object} \ \theta = 0\] so \[W=Fd\] Since you know the only force acting on the object is the force of friction, then \[F_{friction} = \frac{W}{d}\] \[=\frac{-mgh}{d}= -\frac{mg (3m\sin 30º)}{5m}=-\frac{3mg}{10}\] It's negative meaning it opposes the direction of motion. Then using that knowledge, and using the magnitude of the force, the force of friction is defined to be \[F_f = \mu N \rightarrow \text{N is the normal force of the box; N=mg}\] so \[\mu = \frac{F_f}{N} = \frac{ \frac{3mg}{10}}{mg} = .3\] You could do this similarly by using kinematics to find the acceleration of the box using the velocity from the first part. \[v^2 = v_i^2+2ad\] \[a=-\frac{v_i^2}{2d}\] Then summing the forces on the box, giving \[ \sum F = F_f=F_{net}\] \[ F_f = ma\] \[F_f = -\frac{mv^2}{2d}=-\frac{mv^2}{2} \frac{1}{d}\] which is cool because you can see the kinetic energy term just falls out! And we know that mv^2\2 = mgh, proving the other method is also correct! ^_^

OpenStudy (anonymous):

49) \[k=105 \ N/m \\ m=2kg \\ x=.100m \\ d = .25m\] So, again, just like douglaswinslow cooper said, you know that the energy imparted to the spring is \[U_{spring}=\frac{1}{2}kx^2\] and then the kinetic translation (although you don't need to solve for v, since you can just plug in Uspring) is \[E_i=E_f\] \[U_{spring} = K\] \[\frac{1}{2}kx^2= \frac{1}{2}mv^2\] and you know that the work done by friction negates all of the energy given to it by the spring \[W=\Delta K = -\frac{1}{2}mv^2=-\frac{1}{2}kx^2\] \[W=Fd\cos\theta = F_fd\] \[-\frac{1}{2}kx^2=F_fd\] \[F_f = -\frac{kx^2}{2d}\] Then to find the coefficient of friction \[\mu N=F_f\] \[\mu N = \frac{kx^2}{2d}\] \[ \mu = \frac{kx^2}{2dN}\] \[=\frac{kx^2}{2dmg}\] \[ = \frac{(105N/m)(.1m)^2}{2(.25m)(2kg)(9.81m/s^2)}\]

OpenStudy (kittiwitti1):

okay, I need a few days to go over this because I still have ongoing homework (as well as a quiz in 2 days)...

OpenStudy (anonymous):

cool cool ^^